# SoK: Exploring the Potential of Large Language Models for Improving Digital Forensic Investigation Efficiency

### v1

@@ 5.8. Challenges for LLMs in Digital Forensics
To optimise results, LLMs likely need to be trained with specific forensic data (i.e., previous case data) to achieve the best results. Given the complexity and variation of cases, it is questionable how good the training data is and whether there is sufficient data (Breitinger et al., 2024). Biases present in training data can lead to skewed interpretations and unjust outcomes.
@@
Investigative findings and drawing conclusions are also heavily based on investigator experience and intuition, which are difficult to replicate using LLMs. Initially, they will only prove effective for certain subtasks within digital forensics, which raises questions about their superiority over deterministic programs in certain contexts. For instance, data may be parsed/converted using a regular piece of software or an LLM. While the former works deterministically, the latter is more general, but requires the practitioners to validate the output. This leads to the fundamental necessity for explainability (i.e., not using blackboxes) for investigations. The explainability of LLM-generated results remains a daunting challenge, as understanding the reasons for their, frequently non-deterministic, results is often elusive (Michelet and Breitinger, 2023).
@@
Maintaining the energy- and resource-intensive infrastructure required to train and deploy LLMs is a significant financial burden, making it impractical for smaller forensic laboratories to utilise such technologies. Centralised systems may be an option but require clear guidelines on how data can/may be shared and exchanged. While LLMs can serve as valuable tools to support forensic investigations, it must be recognised that they currently function best as an aid, not a substitute for human expertise (Scanlon et al., 2023b). There is a risk that people may place too much trust in the results generated by LLMs (over-reliance), which could lead to complacency and overlook the need for detailed human expert analysis and validation.
@@
Finally, ethical and legal considerations must also be discussed. Determining accountability in cases where LLMs produce false information or are compromised by hacking. Clarifying responsibilities between developers, users and regulators is crucial to establish a framework for accountability. If LLM-generated DF results lead to incorrect information, the responsibility may lie with the developers for ensuring the accuracy of the model and with the users for the appropriate interpretation and validation of the results.
5.9. Risks of Integration
@@
The integration of LLMs within the DF process comes with inherent risks – extra to those general LLM limitations outlined in Section 4.6. Notably, in the examination, analysis, and reporting phases, the utilisation of LLMs introduces the risk of producing inaccurate information, primarily due to the phenomenon of inheritance hallucinations associated with these models (Michelet and Breitinger, 2023; Scanlon et al., 2023b). Additionally, the biases and obscurities present in an inheritance model may significantly impact the performance of a DF-focused LLM – potentially leading to the unacceptable generation of biased or inaccurate information within the DF process.
@@
It is also crucial to acknowledge that DF LLMs, like any complex models, are susceptible to adversarial manipulation (Zou et al., 2023). This vulnerability poses a substantial risk in the context of sensitive domains such as DF, where the integrity of information gained is paramount. Adversarial attacks can compromise the reliability of LLM-generated outputs, potentially influencing the outcomes of various phases within the DF process.
@@
6. Conclusion
@@
The convergence of LLMs with an array of technologies represents an exciting synergy. Although the utilisation of LLMs in the realm of DF is still in its nascent stages, there is evidence of their substantial potential to significantly augment the efficiency of investigations. The exploration of investment for LLMs across the entire DF process is considered, aiming to enhance the productivity and efficiency of investigations. Additionally, the integration of LLMs into current DF tools is posited to reduce user training times, as these models comprehend natural language input and provide outputs accordingly. In the dynamic landscape of LLM applications within DF, promising avenues for further exploration and advancement unfold.
@@
While the surge in LLM research is promising, it is crucial to balance enthusiasm with an awareness of existing challenges. The propensity for LLMs to produce hallucinations highlights the need for human oversight in critical decision-making processes, underscoring the irreplaceable value of human judgement, intuition and expertise. A notable limitation is the language dependency issue, as most LLMs are predominantly trained on English data, reducing their effectiveness with non-English content. Additionally, the deployment of LLMs in DF involves significant costs related to infrastructure for processing evidence. Questions also arise about the validation of task correctness and quality when automated by LLMs, as well as the legal and professional acceptance of results obtained with limited human intervention.
@@
Integrating LLMs with automated agents offers a promising path to automating DF processes, potentially allowing multiple cases to be handled concurrently for more timely and precise outcomes. This integration could significantly streamline investigations. Future research should explore the role of LLMs and AI in DF decision-making. It is essential to focus on validating LLM-generated outputs to ensure their scope, accuracy, reliability, and trustworthiness in investigations. Further studies comparing DF outcomes with and without LLM integration are critical, as they could highlight LLMs’ benefits and controlled applicability in DF and similar fields.
@@
In essence, while LLMs offer exciting prospects for the future of digital forensics, a balanced approach that integrates their strengths with human oversight is essential for harnessing their full potential. Inevitably, LLM-facilitated DF processes will themselves become the focus of future investigation.

### v2

@@ 5.1 Challenges for LLMs in Digital Forensics
To optimise results, LLMs likely need to be trained with specific forensic data (i.e., previous case data) to achieve the best results. Given the complexity and variation of the cases, it is questionable how good the training data are and whether there are sufficient data [158]. Any bias in the training data can lead to skewed interpretations and unjust outcomes.
@@
Investigative findings and conclusions are also highly based on investigator experience and intuition, which are difficult to replicate using LLMs. Initially, they will only be effective for certain subtasks within digital forensics, raising questions about their superiority over deterministic programs in certain contexts. For instance, data may be parsed/converted using regular software or an LLM. While the former works deterministically, the latter is more general, but requires the practitioners to validate the output. This leads to the fundamental necessity for explainability (i.e., not using blackboxes) for investigations. The explainability of LLM-generated results remains a daunting challenge, as understanding the reasons for their, often non-deterministic, results is often elusive [157].
@@
Maintaining the energy and resource intensive infrastructure required to train and deploy LLMs is a significant financial burden, making it impractical for smaller forensic laboratories to utilise such technologies. Centralised systems may be an option, but require clear guidelines on how data can/may be shared and exchanged. Although LLMs can serve as valuable tools to support forensic investigations, it should be recognised that they currently function best as an aid, not a substitute for human expertise [159]. There is a risk that people may place too much trust in the results generated by LLMs (over-reliance), which could lead to complacency and overlook the need for detailed human expert analysis and validation.
To mitigate the potential misuse of LLMs, many LLMs undergo censorship [160, 58]. While this censorship may serve as a preventive measure against unethical use, it can pose challenges in the field of DF. For instance, if an investigator seeks evidence related to ‘drugs’ or evidence of other illegal material, censorship of the LLM may restrict access to accurate information pertaining to the investigator’s query. This limitation underscores the need for a nuanced approach to censorship in LLMs, balancing ethical considerations with the imperative of facilitating effective forensic investigations. In addition, the censorship of generic, publicly accessible LLMs further supports the argument for a discipline specific DF LLM.
@@
Finally, ethical and legal considerations must also be discussed. Determining accountability in cases where LLMs produce false information or are compromised by hacking. Clarifying responsibilities between developers, users, and regulators is crucial to establish a framework for accountability. If LLM generated digital forensics (DF) results lead to incorrect information, the responsibility may lie with both the developers, for ensuring the model’s accuracy, and the users, for appropriately interpreting and validating the results.
5.2 Risks of Integration
@@
The integration of LLMs within the DF process comes with inherent risks – extra to those general LLM limitations outlined in Section 3.7. Notably, in the examination, analysis, and reporting phases, the utilisation of LLMs introduces the risk of producing inaccurate information, primarily due to the phenomenon of inheritance hallucinations associated with these models [157, 159]. Additionally, the biases and obscurities present in an inheritance model can significantly impact the performance of a DF-focused LLM – potentially leading to the unacceptable generation of biased or inaccurate information within the DF process.
@@
It is also crucial to acknowledge that DF LLMs, like any complex model, are susceptible to adversarial manipulation [161]. This vulnerability poses a substantial risk in the context of sensitive domains such as DF, where the integrity of the information obtained is paramount. Adversarial attacks can compromise the reliability of LLM-generated outputs, potentially influencing the outcomes of various phases within the DF process.
@@
Mitigating these changes and risks can be challenging, particularly in scenarios involving adopting country-specific legal systems. However, there are potential strategies to address technical challenges such as hallucination, censorship, and substantial infrastructure costs.
One solution to mitigate hallucination was suggested by Ji et al. [162], who proposed an interactive self-reflection method for generated knowledge and answers – an approach that has been demonstrated promise. Another hallucination mitigation method is using RAG, which provides a larger knowledge base to the LLMs to minimise unknown information [132]. Other methods such as knowledge graphs, bias detection mechanisms, active learning methods for LLMs, supervised fine-tuning strategies, hallucination mitigation frameworks and new decoding strategies can also help mitigate hallucination to some extent [163, 164].
Censorship issues can be addressed by fine-tuning the model with uncensored information, a technique already applied to the LLaMA and Mistral models, leading to the development of the Dolphin models. An example is the Dolphin-2.0-mistral-7b, which is an uncensored version of the Mistral 7B model [165].
The high infrastructure costs associated with these models can be mitigated by employing Data Forensics as a Service (DFaaS) platforms such as Hansken. With DFaaS, investigators need only input queries related to their investigations using personal computers, while the platform manages the model maintenance and computational demands [29].
6 Conclusion
@@
The convergence of LLMs with an array of technologies represents an exciting synergy. Although the utilisation of LLMs in the realm of DF is still in its nascent stages, there is evidence of their substantial potential to significantly augment the efficiency of investigations. The exploration of investments for LLMs across the entire DF process is considered, with the aim of improving the productivity and efficiency of investigations. Additionally, the integration of LLMs into current DF tools is posited to reduce user training times, as these models comprehend natural language input and provide output accordingly. In the dynamic landscape of LLM applications for DF, promising avenues for further exploration and advancement unfold.
@@
While the surge in LLM research is promising, it is crucial to balance enthusiasm with an awareness of existing challenges. The propensity for LLMs to produce hallucinations highlights the need for human oversight in critical decision-making processes, underscoring the irreplaceable value of human judgement, intuition, and expertise. A notable limitation is the language dependency issue, as most LLMs are predominantly trained on English data, reducing their effectiveness with non-English content. Additionally, the deployment of LLMs in DF involves significant costs related to infrastructure for processing evidence. Questions also arise about the validation of task correctness and quality when automated by LLMs, as well as the legal and professional acceptance of results obtained with limited human intervention.
The trustworthiness of LLMs remains a debatable issue that requires careful attention. It is crucial to establish clear boundaries and measures to define LLM trustworthiness. Addressing this will be a key aspect in the field of DF, ensuring that LLMs can be relied upon for accurate and secure analysis, with explainability of their operations being paramount.
@@
Integrating LLMs with automated agents offers a promising path to automating DF processes, potentially allowing multiple cases to be handled concurrently for more timely and precise outcomes. This integration could significantly streamline investigations. Future research should explore the role of LLMs and AI in DF decision-making. It is essential to focus on validating LLM generated outputs to ensure their scope, accuracy, reliability, and trustworthiness in investigations. Further studies comparing DF outcomes with and without LLM integration are critical, as they could highlight LLMs’ benefits and controlled applicability in DF and similar fields.
@@
In essence, while LLMs offer exciting prospects for the future of digital forensics, a balanced approach that integrates their strengths with human oversight is essential for harnessing their full potential. Inevitably, LLM facilitated DF processes will themselves become the focus of future investigation.